How many milliliters of a 5% acid solution and how many milliliters of a 17% acid solution must be mixed to obtain 60 mL of a 13% acid solution?
Found 2 solutions by josgarithmetic, addingup:
Answer by josgarithmetic(39617)
Answer by addingup(3677)

0.05x+0.17(60-x) = 0.13(60)
0.05x+-0.17x+10.2 = 7.8
-0.12x = -2.4
x = 20 you need 20ml of the 5% solution and 60-20 = 40ml of the 17% solution.